CuO can be formed from the reaction of 18.7g of CuS and 12.0g of O2. Which reactant will run out first


1
Expert's answer
2021-02-28T06:19:33-0500

Solution:

The balanced chemical equation:

2CuS + 3O2 → 2CuO + 2SO2

According to the equation above: n(CuS)/2 = n(O2)/3


Determine moles of each reactant:

18.7 g CuS × (1 mol CuS / 95.611 g CuS) = 0.1956 mol CuS

12.0 g O2 × (1 mol O2 / 31.998 g O2) = 0.3750 mol O2


Choose one reactant and determine how many moles of the other reactant are necessary to completely react with it. Let's choose CuS:

n(O2) = 3 × n(CuS) / 2 = (3 × 0.1956 mol) / 2 = 0.2934 mol


The calculation above means that we need 0.2934 mol of O2 to completely react with CuS.

We have 0.3750 mol of O2 and therefore more than enough oxygen.

Thus oxygen (O2) is in excess and copper monosulfide (CuS) must be the limiting reagent.

Hence, CuS will run out first.


Answer: CuS will run out first.